### v3

@@ 1 INTRODUCTION
Especially, the practical value of PLD emerges in scenarios where adversaries can suffer from inaccurate information. For example, in military tactical communications, forward units can transmit target coordinates as deceptively-encoded messages: intercepting adversaries decode falsified coordinates that misdirect their resources, while legitimate receivers reconstruct true coordinates. In confidential business communications, as another instance, corporate executives can transmit merger target identifiers or product launch schedules in the same way: malicious insiders or competitors with leaked protocols intercept falsified information that triggers counterproductive actions, while authorized partners receive true intelligence. Such active deception proves more damaging to adversaries than signal denial, as they commit resources based on false intelligence rather than merely lacking information. We have later proven that PLD is a semantic encryption solution for secure transmission, which relies on the secrecy of the semantic knowledge instead of that of the messages [5].
As the core part of the PLD framework, the deceptive ciphering algorithm is required to fulfill two specific requirements. First, the codebooks of ciphertext and plaintext have to share a same codewrd set, i.e. every possible ciphertext codeword must also be a valid plaintext codeword. Second, every individual codeword must contain meaningful semantic information. While the preliminary works on PLD have been only focusing on the link-level design, little has been discussed about the application scenario or the system-level design, which can be constrained by the above requirements. essential semantic encoding mechanisms to generate meaningful semantic codewords. This creates a critical implementation gap between theoretical PLD concepts and practical semantic communication systems.
@@
Semantic encryption Irrelevant Neccesary, unspecified AI-driven visual encryption
@@ 3 THE VENENA FRAMEWORK
We consider an wiretap channel model with a sender (Alice), a legitimate receiver (Bob), and an eavesdropper (Eve). Alice wants to transmit a short confidential message (e.g., a number) to Bob, while preventing Eve from correctly recovering the message.
@@ III-B DECEPTIVE VISUAL ENCRYPTION
To enhance the security of the visual encryption scheme, we propose a novel framework called VENENA, which leverages the concept of PLD and image poisoning. As shown in Fig.1(b), instead of simply sending an image with the message-associated semantic tag, Alice first randomly generates a falsified message, which is then visual-encoded into a correspondingly tagged image, which we call the original image. This image serves as a deceptive message, since it carries deceptive semantics. It is then processed with a poisoning network in such a way, that the poisoned image at output shall be classified with the true semantic tag, while remaining miminally diversed from the original image in the sense of error power. It is this poisoned image that carries the true semantics. Afterwards, a poison mask is obtained by pixel-wise differentiating the original image from the poisoned one. Seen from the semantic PLD perspective, the poison mask carries a random knowlege that maintains the confidentiality of the true semantics, when both the semantic codec and the deceptive message are leaked [5]. The original image and the poison mask are then multiplexed and sent together, supposed to be decoded by Bob separately for construction of the poisoned image. Specifically, while VENENA is agnostic to the multiplexing method, in this work we apply a power-domain non-orthogonal multiplexing, where the original image is allocated with significantly higher power than the poison mask, so that it can be first decoded treating the poison mask as a weak inteference. Subsequently, the decoded original image is subtracted from the received signal, and the poison mask can be decoded from the residual signal (known as the technique of successive interference cancellation (SIC) [15]). The power allocation shall be fine tuned in such way, that the signal-to-interference-and-noise ratio (SINR) at Bob (i) significantly exceeds the necessary level for successful decoding of the original image, but (ii) fulfills the requirement for decoding of the poison mask with only a small margin. Thus, PLS is selectively applied on the poison mask component, which carries the differential information between the true and falsified messages, while the original image carrying falsified message is well exposed to potential eavesdroppers.
@@ V-A SIMULATION SETUP
Shall VENENA be applied in scenarios where high data rates are required, we recommend to rely on offline-poisoned image databases instead of carrying out poisoning online. With poisoned images pre-generated and stored for immediate use during transmission, reducing the time complexity from $O(n_{\text{iter}}\cdot L\cdot d)$ to $O(1)$ per image. An occoasional update of the poisoned image database shall be scheduled in this case, to eliminate long-term statistical learning by eavesdroppers.
@@
    Full knowledge, where Eve shares the identical knowledge with Bob about the system design, including the VENENA framework, the protocol, the data format, the parameter specifications, the image classification network used by Bob, as well as the mapping table; and
@@
In this article, we focus on the case where Eve to be consistently eavesdropping. For readers who are interested in sophisticated dynamic eavesdropping strategies, we refer to our related work [5].

### v4

@@ 1 INTRODUCTION
Especially, the practical value of PLD emerges in scenarios where adversaries can suffer from inaccurate information. For example, in military tactical communications, forward units can transmit target coordinates as deceptively-encoded messages: intercepting adversaries decode falsified coordinates that misdirect their resources, while legitimate receivers reconstruct true coordinates. In confidential business communications, as another instance, corporate executives can transmit merger target identifiers or product launch schedules in the same way: malicious insiders or competitors with leaked protocols intercept falsified information that triggers counterproductive actions, while authorized partners receive true intelligence. Such active deception proves more damaging to adversaries than signal denial, as they commit resources based on false intelligence rather than merely lacking information. We have later proven that PLD is a semantic encryption solution for secure transmission, which relies on the secrecy of the semantic knowledge instead of that of the messages [5, 6]. Compared to contemporary works in adversarial semantic communication that focus on the attack-oriented generation of semantic perturbations–such as SemAdv [7], SemBLK [8], and SemPerGe [9]–or on protecting semantic communication systems against such perturbations–such as SemProtector [10]PLD targets a fundamentally different objective. Specifically, rather than suppressing semantic distortion, PLD discriminatively induces semantic distortions at illegitimate eavesdroppers while preserving semantically correct perception at the legitimate receiver.
As the core part of the PLD framework, the deceptive ciphering algorithm is required to fulfill two specific requirements. First, the codebooks of ciphertext and plaintext have to share a same codewrd set, i.e. every possible ciphertext codeword must also be a valid plaintext codeword. Second, every individual codeword must contain meaningful semantic information. While the preliminary works on PLD [2, 3, 4, 5, 6] have provided in-depth information-theoretic and semantic analyses–on perspectives of deception rate, leakage-failure rate, and semantic distortion–as the cornerstone, they have been only focusing on the link-level design, little has been discussed about the application scenario or the system-level design, which can be constrained by the above requirements. essential semantic encoding mechanisms to generate meaningful semantic codewords. This creates a critical implementation gap between theoretical PLD concepts and practical semantic communication systems.
@@
Semantic encryption Irrelevant Necessary, unspecified AI-driven visual encryption
@@ 3 THE VENENA FRAMEWORK
We consider an wiretap channel model with a sender (Alice), a legitimate receiver (Bob), and an eavesdropper (Eve). Alice wants to transmit a short confidential message (e.g., a number) to Bob, while preventing Eve from correctly recovering the message. We consider Eve as able to continuously monitor the wireless channel and intercept all transmissions from Alice to Bob. It may have or not have full knowledge about the details of the system design. Both the legitimate and eavesdropping channels are lossy, yet with sufficiently high capacity to transmit an image at required data rate.
@@ III-B DECEPTIVE VISUAL ENCRYPTION
To enhance the security of the visual encryption scheme, we propose a novel framework called VENENA, which leverages the concept of PLD and image poisoning. As shown in Fig.1(b), instead of simply sending an image with the message-associated semantic tag, Alice first randomly generates a falsified message, which is then visual-encoded into a correspondingly tagged image, which we call the original image. This image serves as a deceptive message, since it carries deceptive semantics. It is then processed with a poisoning network in such a way, that the poisoned image at output shall be classified with the true semantic tag, while remaining miminally diversed from the original image in the sense of error power. It is this poisoned image that carries the true semantics. Afterwards, a poison mask is obtained by pixel-wise differentiating the original image from the poisoned one. Seen from the semantic PLD perspective, the poison mask carries a random knowledge that maintains the confidentiality of the true semantics, when both the semantic codec and the deceptive message are leaked [5]. The original image and the poison mask are then multiplexed and sent together, supposed to be decoded by Bob separately for construction of the poisoned image. Specifically, while VENENA is agnostic to the multiplexing method, in this work we apply a power-domain non-orthogonal multiplexing, where the original image is allocated with significantly higher power than the poison mask, so that it can be first decoded treating the poison mask as a weak inteference. This introduction of inteference, of course, will reduce the signal-to-interference-and-noise ratio (SINR) of the received original image and therewith the classification accuracy for both Bob and Eve. Subsequently, the decoded original image is subtracted from the received signal, and the poison mask can be decoded from the residual signal (known as the technique of successive interference cancellation (SIC) [20]). The power allocation shall be fine tuned in such way, that the SINR at Bob (i) significantly exceeds the necessary level for successful decoding of the original image, but (ii) fulfills the requirement for decoding of the poison mask with only a small margin. Thus, PLS is selectively applied on the poison mask component, which carries the differential information between the true and falsified messages, while the original image carrying falsified message is well exposed to potential eavesdroppers.
@@ V-A SIMULATION SETUP
Shall VENENA be applied in scenarios where high data rates are required, we recommend to rely on offline-poisoned image databases instead of carrying out poisoning online. With poisoned images pre-generated and stored for immediate use during transmission, reducing the time complexity from $O(n_{\text{iter}}\cdot L\cdot d)$ to $O(1)$ per image. An occasional update of the poisoned image database shall be scheduled in this case, to eliminate long-term statistical learning by eavesdroppers.
@@
    Full knowledge, where Eve shares the identical knowledge with Bob about the system design, including the VENENA framework, the protocol, the data format, the parameter specifications, the image classification network used by Bob, as well as the mapping table. In this case, Eve will try to carry out SIC to decode both the original image and the poison mask, and then reconstruct the poisoned image for message perception. If it fails to decode the poison mask, it will assume that VENENA is not activated, and perceive the message carried by the original image only.
@@ VI-A ENCRYPTION CODING EFFICIENCY
The specific implementation of these approaches, however, are out of the scope of this article, and shall be left for future works.
@@ VI-C CHANNEL CONDITION ASSUMPTIONS
VI-D ADVANCED EAVESROPPERS
In this work, we have considered an informed Eve with full knowledge of the system design, and verified VENENA’s secrecy thereagainst. However, even more sophisticated eavesdropping strategies could be envisioned. For instance, Eve can adapt its receiption and decryption strategy dynamically according to real-time channel conditions, which we have preliminarily explored in a generic and theoretical context of semantic PLD in [5], and recently discussed in more depth in [6]. Besides, multiple coordinated eavesdroppers may share their received signals and therewith overwhelm Bob regarding the SINR, in order to violate the secrecy. These advanced eavesdropping strategies pose new challenges to VENENA, motivating future research on more resilient encryption strategies, which are out of the scope of this article.